Select the x-coordinate of the vertex of the parabola defined by the function f(x) = -7x2 + 3x + 5.
Answer by CubeyThePenguin(3113)

-7x^2 + 3x + 5

vertex at x = -b/2a = -3/(2 * -7) = 3/14

f(3/14) = 149/28

vertex: (3/14, 149/28)